# Galaxy Zoo: Cosmic Dawn – morphological classifications for over 41 000 galaxies in the Euclid Deep Field North from the Hawaii Two-0 Cosmic Dawn survey

### v1

@@ Abstract
We present morphological classifications of over 41 000 galaxies out to $z_{\rm phot}\sim 2.5$ across six square degrees of the Euclid Deep Field North (EDFN) from the Hawaii Twenty Square Degree (H20) survey, a part of the wider Cosmic Dawn survey. Galaxy Zoo citizen scientists play a crucial role in the examination of large astronomical data sets through crowdsourced data mining of extragalactic imaging. This iteration, Galaxy Zoo: Cosmic Dawn (GZCD), saw tens of thousands of volunteers and the deep learning foundation model Zoobot collectively classify objects in ultra-deep multiband Hyper Suprime-Cam (HSC) imaging down to a depth of $m_{HSC-i}=21.5$. Here, we present the details and general analysis of this iteration, including the use of Zoobot in an active learning cycle to improve both model performance and volunteer experience, as well as the discovery of 51 new gravitational lenses in the EDFN. We also announce the public data release of the classifications for over 45 000 subjects, including more than 41 000 galaxies (median $z_{\rm phot}$ of $0.42\pm 0.23$), along with their associated image cutouts. This data set provides a valuable opportunity for follow-up imaging of objects in the EDFN as well as acting as a truth set for training deep learning models for application to ground-based surveys like that of the newly operational Vera C. Rubin Observatory.
@@
pubyear: 2025pagerange: Galaxy Zoo: Cosmic Dawn – morphological classifications for over 41 000 galaxies in the Euclid Deep Field North from the Hawaii Two-0 Cosmic Dawn surveyA
@@ 1 Introduction
The visual morphology of galaxies provides a valuable means of studying their structure and evolution over cosmic time. Properties derived from morphology pertain to a galaxy’s history, such as the development of discs and spiral structure (Masters et al., 2019; Lingard et al., 2021; Mengistu & Masters, 2023), the assembly of bulges and bars and the role they play in quenching (Lin et al., 2019; Géron et al., 2021, 2024), the activity of its central black hole (Garland et al., 2024), and any previous or current merging events (Martin et al., 2018). Such morphological properties can be cross-correlated with one another and compared to, for example, the rate of star formation within galaxies (Willett et al., 2015; Hart et al., 2017; Porter-Temple et al., 2022). Examining morphology can also reveal rare objects, such as low surface brightness (LSB) galaxies which are inherently difficult to detect (Bothun et al., 1997; Impey & Bothun, 1997), or galaxies acting as gravitational lenses due to their dark matter haloes distorting the light of higher redshift background galaxies situated almost perfectly along the line of sight (e.g. Treu, 2010; Lemon et al., 2024; Euclid Collaboration: Walmsley et al., 2025b). With the combination of many properties contributing to a galaxy’s morphology, a large number of classified galaxies are required to study the effect of each of these in detail. The collection of such large samples of galaxies is now underway with the latest and upcoming large-scale surveys like the Euclid survey (Laureijs et al., 2011; Euclid Collaboration: Mellier et al., 2025) and the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST; Ivezić et al., 2019), which are pushing toward deeper imaging at higher redshifts while maintaining wide survey areas, in turn allowing the study of galaxy evolution at earlier epochs.
@@
With these surveys observing anywhere from tens of thousands to many millions of objects, image inspection methods are needed in order to study them in detail. Citizen science, specifically crowdsourced data mining, plays a crucial role in this task, utilising the collective effort of thousands of members of the public (‘volunteers’) to classify large quantities of data much faster than could be achieved by professional researchers alone. Galaxy Zoo (Lintott et al., 2008) is the longest running project on the Zooniverse citizen science online platform, regularly inviting tens of thousands of volunteers to take part in classifying images of galaxies through a web interface. It has seen a number of iterations over its 17-year lifespan involving collaborations with a number of surveys (e.g. Willett et al., 2013; Simmons et al., 2017; Willett et al., 2017; Walmsley et al., 2022a), and their results have resulted in numerous further studies of galaxy morphology (a review is presented in Masters & Galaxy Zoo Team, 2020).
@@
An obvious complement to citizen science in image classification is the use of machine learning, especially with the growing prevalence of deep learning algorithms in astronomy, including galaxy morphology (e.g. Huertas-Company & Lanusse, 2023; Cao et al., 2024; Etsebeth et al., 2024; Mohale & Lochner, 2024; Tohill et al., 2024; Lochner & Rudnick, 2025). Machine learning can be used to automatically classify images almost instantaneously, but for the desired discrete classifications this first requires training the machine learning model on tens to hundreds of thousands of pre-classified images. Additionally, while such models can be used to classify common features like spiral arms and bulge shapes, they often struggle with features that are complex or rare in their training set such as gravitational lenses, irregular galaxies, and mergers. Meanwhile, volunteers have the additional benefit of being able to identify particularly unusual and interesting objects perhaps never seen before through serendipitous discovery. Nevertheless, the far higher classification rate of machine learning makes it desirable as the data sets of images become larger and larger, and so in recent years Galaxy Zoo has employed its own machine learning model, Zoobot (Walmsley et al., 2023a), alongside its crowdsourcing effort (e.g. Walmsley et al., 2020, 2022a, 2022b, 2023b; Euclid Collaboration: Aussel et al., 2024; Euclid Collaboration: Walmsley et al., 2025a). Having trained on classifications from previous Galaxy Zoo iterations, this model has been used for pre-classification of the simplest objects as well as part of an active learning cycle to continuously retrain as new volunteer classifications arrive in order to provide volunteers with a curated sample of the more difficult and often more interesting objects. This combination of citizen science and machine learning thus provides deeper engagement for volunteers while also speeding up the rate of classifications to tackle the growing size of astronomical data sets. Zoobot is now considered to be a foundation model for extragalactic research (Walmsley, 2024; Euclid Collaboration: Walmsley et al., 2025a) and is being employed beyond Galaxy Zoo to other astronomical tasks (e.g. Popp et al., 2024; Euclid Collaboration: Lines et al., 2025; Pearce-Casey et al., 2025).
@@
In this data release, we present results and data sets from one of the latest Galaxy Zoo iterations, Galaxy Zoo: Cosmic Dawn (GZCD), in which Zoobot-assisted volunteers collectively classify tens of thousands of colour postage stamp images from the multiwavelength Cosmic Dawn survey (Euclid Collaboration: McPartland et al., 2025) of the Euclid deep fields. In particular, these ultra-deep multiband Hyper Suprime-Cam (HSC) images come from Cosmic Dawn’s ongoing Hawaii Twenty Square Degree (H20) survey (Euclid Collaboration: Zalesky et al., 2025), covering the first six square degrees of the Euclid Deep Field North (EDFN) in the North Ecliptic Pole (NEP). These deep fields are some of the darkest areas of the sky, selected by Cosmic Dawn for multiple multiwavelength studies in part as preparation for the ultra-deep photometry and spectroscopy of the Euclid mission. As such, classifications made by volunteers can provide a means of selecting the most interesting objects for further examination with Euclid imaging and follow-up observations. For example, the classifications made by volunteers may help improve our understanding of galaxy evolution over a significant fraction of the age of the Universe through identifying rarer objects like galaxy mergers to investigate the role they play in influencing the star-formation rates of galaxies (the extent of how and when this enhancement occurs is still uncertain; Pearson et al., 2019; Feng et al., 2024; Ferreira et al., 2025).
@@
Compared to most previous Galaxy Zoo iterations, this imaging is significantly deeper, which hence allows for the identification of more LSB features, and this iteration also places greater emphasis on searching for (higher-redshift) clumpy galaxies and gravitational lenses (see Section 3.1), as well as extremely red objects (see Section 3.2). HSC is also a precursor for the Legacy Survey of Space and Time (LSST) Vera C. Rubin Observatory, and so the morphological classifications in this data release can be utilised as multiband ground truth sets in training deep learning models for application to Rubin imaging.
@@
GZCD ran on the Zooniverse platform for over six months, with tens of thousands of volunteers providing almost four million classifications. After the first third of the subjects were fully classified, a pre-trained Zoobot was further trained on these and deployed in an active learning cycle with volunteers for the remaining two thirds of the subjects. A final run of Zoobot was also applied to the entire subject set at the end of the project.
@@
This paper is organised as follows. In Section 2 we discuss the imaging data used for this work, detailing the surveys, subject selection & cutout creation methods, and the metadata associated with each subject. In Section 3, further details regarding the project implementation are provided, including the decision tree changes, extra tagging task, the project launch, the use of Zoobot machine learning, and data aggregation at the conclusion of the project. Results are presented in Section 4, including a statistical overview of the various morphological classifications and a comparison between volunteer and Zoobot responses. This is followed by a discussion of the data release catalogues in Section 5 and final conclusions in Section 6.
@@ 2.1 Cosmic Dawn and Hawaii Twenty Square Degree surveys
The Cosmic Dawn survey covers 59 square degrees of the Euclid Deep and Auxiliary Fields from UV/optical to mid-infrared with the aim to understand how galaxies, black holes and dark matter haloes co-evolve from the epoch of reionization to the present – more details of this survey can be found in Euclid Collaboration: McPartland et al. (2025). One facet of this is the ongoing H20 survey, which is covering the two primary Euclid deep calibration fields with $u$-band imaging from the Canada-France-Hawaii Telescope (CFHT) MegaCam instrument and ultra-deep grizgriz imaging from HSC, the latter of which is on board the 8.2-metre Subaru telescope on the summit of Maunakea in Hawaii. H20 aims to push the boundaries of extragalactic astronomy by studying galaxy evolution and co-evolution with dark matter haloes out to a high redshift of $z=7$, through combination with Spitzer Space Telescope infrared imaging and Keck DEIMOS (DEep Imaging Multi-Object Spectrograph) spectroscopy. Targeting the entire twenty square degrees of the EDFN, H20 HSC imaging has intentionally similar depth to Euclid’s own near-infrared imaging in this field, with $5\sigma$ depths of $\sim$27 AB mag (Euclid Collaboration: Zalesky et al., 2025) accompanied by a 0.168 arcsec per pixel angular resolution. HSC imaging is seeing-limited, with a median seeing FWHM of 0.67 arcsec over several years of imaging (Miyazaki et al., 2018); variations in the PSF between 0.4 and 1.0 arcsec are observed for the H20 survey (Figure 4 of Euclid Collaboration: Zalesky et al., 2025).
The subjects used in this iteration of Galaxy Zoo consist of ultra-deep grigri HSC images that cover the first six square degrees of the EDFN observed at full depth by the H20 survey. For reference, the sky coverage of the 9.37 square degrees of the Pre-Launch (PL) EDFN Cosmic Dawn catalogue can be seen in Figure 1 of Euclid Collaboration: Zalesky et al. (2025), and more details of how the H20 data was obtained and processed can be found in that work.
@@ 2.2 Subject selection
The subjects consist of deep multi-band cutout images centred on objects identified by the H20 pipeline, limited to those with effective radii $r_{\rm eff}\geq 0.6$ arcsec as the HSC PSF’s 0.67 arcsec FWHM made smaller objects too difficult to classify morphologically. In rare cases (<<1 per cent of all detected sources), the pipeline failed to produce suitable photometric models due to a bright neighbouring object, and so for these sources and point sources the detection-measured major axis was used as a proxy for the effective radius. By taking into account assumptions regarding the average background noise, estimates of the HSC-$i$ band magnitudes for these subjects were obtained that were accurate to within $\sim$1 mag, for use in creating image cutouts (see Section 2.3).
@@
Figure 1: Distribution of photometric redshifts for the subjects in GZCD as measured by the H20 pipeline. Subjects for which the pipeline failed to produce suitable photometric redshift estimates are not included. These distributions cover the subject set so stars, artifacts and ‘bad zoom’ subjects have not been excluded.
Figure 2: Distribution of HSC-$g$, -$r$, and -$i$ band apparent magnitudes for the subjects in GZCD as measured by the H20 pipeline. Subjects for which the pipeline failed to produce suitable photometric models are not included. These distributions cover the subject set so stars, artifacts and ‘bad zoom’ subjects have not been excluded.
The photometric redshift distribution for this subject set can be seen in Figure 1: while 95.4 per cent have $0<z_{\rm phot}<1.0$ (median of $0.42\pm 0.23$), the remaining 4.6 per cent extend up to $z_{\rm phot}<=6.7$. However, it should be noted that for 19.5 per cent of subjects, photometric redshift estimation from the H20 pipeline failed and so these subjects are not included in this range. As discussed in Section 4.1.2, the analysis in later sections is limited to the 45 742 subjects with $z_{\rm phot}<2.5$. Figure 2 shows the distribution of HSC-$g$, -$r$, and -$i$ band apparent magnitudes for the subject set, excluding the small fraction of those for which the modelling failed. Ranges for these magnitudes are $5.64<m_{{\rm HSC}-g}<28.81$, $4.95<m_{{\rm HSC}-r}<29.30$, and $5.22<m_{{\rm HSC}-i}<21.5$, with median averages of 22.03, 21.07, and 20.55, respectively.

### v2

@@ Abstract
We present morphological classifications of over 41 000 galaxies out to $z_{\rm phot}\sim 2.5$ across six square degrees of the Euclid Deep Field North (EDFN) from the Hawaii Twenty Square Degree (H20) survey, a part of the wider Cosmic Dawn survey. Galaxy Zoo citizen scientists play a crucial role in the examination of large astronomical data sets through crowdsourced data mining of extragalactic imaging. This iteration, Galaxy Zoo: Cosmic Dawn (GZCD), saw tens of thousands of volunteers and the deep learning foundation model Zoobot collectively classify objects in ultra-deep multiband Hyper Suprime-Cam (HSC) imaging down to a depth of $m_{HSC-i}=21.5$. Here, we present the details and general analysis of this iteration, including the use of Zoobot in an active learning cycle to improve both model performance and volunteer experience, as well as the discovery of 51 new gravitational lenses in the EDFN. We also announce the public data release of the classifications for over 45 000 subjects, including more than 41 000 galaxies (median $z_{\rm phot}$ of $0.42\pm 0.23$), along with their associated image cutouts. This data set provides a valuable opportunity for follow-up imaging of objects in the EDFN as well as acting as a truth set for training deep learning models for application to ground-based surveys like that of the Ultraviolet Near-Infrared Optical Northern Survey (UNIONS) collaboration and the newly operational Vera C. Rubin Observatory.
@@
pubyear: 2026pagerange: Galaxy Zoo: Cosmic Dawn – morphological classifications for over 41 000 galaxies in the Euclid Deep Field North from the Hawaii Two-0 Cosmic Dawn surveyA
@@ 1 Introduction
The visual morphology of galaxies provides a valuable means of studying their structure and evolution over cosmic time. Properties derived from morphology pertain to a galaxy’s history, such as the development of discs and spiral structure (Masters et al., 2019; Lingard et al., 2021; Mengistu & Masters, 2023), the assembly of bulges and bars and the role they play in quenching (Lin et al., 2019; Géron et al., 2021, 2024), the activity of its central black hole (Garland et al., 2024), and any previous or current merging events (Martin et al., 2018). Such morphological properties can be cross-correlated with one another and compared to, for example, the rate of star formation within galaxies (Willett et al., 2015; Hart et al., 2017; Porter-Temple et al., 2022). Examining morphology can also reveal rare objects, such as low surface brightness (LSB) galaxies which are inherently difficult to detect (Bothun et al., 1997; Impey & Bothun, 1997), or galaxies acting as gravitational lenses due to their dark matter haloes distorting the light of higher redshift background galaxies situated almost perfectly along the line of sight (e.g. Treu, 2010; Lemon et al., 2024; Euclid Collaboration, 2025b).
With the combination of many properties contributing to a galaxy’s morphology, a large number of classified galaxies are required to study the effect of each of these in detail. The collection of such large samples of galaxies is now underway with the latest and upcoming large-scale surveys like the Euclid survey (Laureijs et al., 2011; Euclid Collaboration, 2025f) and the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST; Ivezić et al., 2019), which are pushing toward deeper imaging at higher redshifts while maintaining wide survey areas, in turn allowing the study of galaxy evolution at earlier epochs. While the Euclid Wide Survey will cover 14 000 square degrees (more than a third of the sky) in both the northern and southern hemispheres, the Rubin LSST Wide Fast Deep (WFD) Survey will cover 18 000 square degrees of the southern sky repeatedly over 10 years. Additionally, the Ultraviolet Near-Infrared Optical Northern Survey (UNIONS; Gwyn et al., 2025) has been obtaining imaging from the combination of the Canada-France-Hawaii Telescope (CFHT), Pan-STARRS, and Subaru Hyper Suprime-Cam (HSC) across 6250 square degrees of the northern sky, acting as a northern complement to Rubin LSST while overlapping with the sky coverage of Euclid.
@@
With these surveys observing anywhere from tens of thousands to many millions of objects, image inspection methods are needed in order to study them in detail. Citizen science, specifically crowdsourced data mining, plays a crucial role in this task, utilising the collective effort of thousands of members of the public (‘volunteers’) to classify large quantities of data much faster than could be achieved by professional researchers alone. Galaxy Zoo (Lintott et al., 2008) is the longest running project on the Zooniverse citizen science online platform, regularly inviting tens of thousands of volunteers to take part in classifying images of galaxies through a web interface. It has seen many iterations over its 17-year lifespan involving collaborations with a number of surveys (e.g. Willett et al., 2013; Simmons et al., 2017; Willett et al., 2017; Walmsley et al., 2022a), and their results have resulted in numerous further studies of galaxy morphology (a review is presented in Masters & Galaxy Zoo Team, 2020).
@@
An obvious complement to citizen science in image classification is the use of machine learning, especially with the growing prevalence of deep learning algorithms in astronomy, including galaxy morphology (e.g. Huertas-Company & Lanusse, 2023; Cao et al., 2024; Etsebeth et al., 2024; Mohale & Lochner, 2024; Tohill et al., 2024; Lochner & Rudnick, 2025). Machine learning can be used to automatically classify images almost instantaneously, but for the desired discrete classifications this first requires training the machine learning model on tens to hundreds of thousands of pre-classified images. Additionally, while such models can be used to classify common features like spiral arms and bulge shapes, they often struggle with features that are complex or rare in their training set such as gravitational lenses, irregular galaxies, and mergers. Meanwhile, volunteers have the additional benefit of being able to identify particularly unusual and interesting objects perhaps never seen before through serendipitous discovery. Nevertheless, the far higher classification rate of machine learning makes it desirable as the data sets of images become larger and larger, and so in recent years Galaxy Zoo has employed its own machine learning model, Zoobot (Walmsley et al., 2023a), alongside its crowdsourcing effort (e.g. Walmsley et al., 2020, 2022a, 2022b, 2023b; Euclid Collaboration, 2024, 2025a). Having trained on classifications from previous Galaxy Zoo iterations, this model has been used for pre-classification of the simplest objects as well as part of an active learning cycle to continuously retrain as new volunteer classifications arrive in order to provide volunteers with a curated sample of the more difficult and often more interesting objects. This combination of citizen science and machine learning thus provides deeper engagement for volunteers while also speeding up the rate of classifications to tackle the growing size of astronomical data sets. Zoobot is now considered to be a foundation model for extragalactic research (Walmsley, 2024; Euclid Collaboration, 2025a) and is being employed beyond Galaxy Zoo to other astronomical tasks (e.g. Popp et al., 2024; Euclid Collaboration, 2025c; Pearce-Casey et al., 2025).
@@
In this data release, we present results and data sets from one of the latest Galaxy Zoo iterations, Galaxy Zoo: Cosmic Dawn (GZCD), in which Zoobot-assisted volunteers collectively classify tens of thousands of colour postage stamp images from the multiwavelength Cosmic Dawn survey (Euclid Collaboration, 2025e) of the Euclid deep fields. In particular, these ultra-deep multiband HSC images come from Cosmic Dawn’s ongoing Hawaii Twenty Square Degree (H20) survey (Euclid Collaboration, 2025d), covering the first six square degrees of the Euclid Deep Field North (EDFN) in the North Ecliptic Pole (NEP). These deep fields are some of the darkest areas of the sky, selected by Cosmic Dawn for multiple multiwavelength studies in part as preparation for the ultra-deep photometry and spectroscopy of the Euclid mission. As such, classifications made by volunteers can provide a means of selecting the most interesting objects for further examination with Euclid imaging and follow-up observations. For example, the classifications may help improve our understanding of galaxy evolution over a significant fraction of the age of the Universe through identifying rarer objects like galaxy mergers to investigate the role they play in influencing the star-formation rates of galaxies (the extent of how and when this enhancement occurs is still uncertain; Pearson et al., 2019; Feng et al., 2024; Ferreira et al., 2025).
@@
Compared to most previous Galaxy Zoo iterations, this imaging is significantly deeper, which hence allows for the identification of more LSB features, and this iteration also places greater emphasis on searching for (higher-redshift) clumpy galaxies and gravitational lenses (see Section 3.1), as well as extremely red objects (see Section 3.2). HSC is also a precursor for Rubin LSST, and so the morphological classifications in this data release can be utilised as multiband ground truth sets in training deep learning models for application to both Rubin imaging and the HSC imaging of UNIONS.
@@
GZCD ran on the Zooniverse platform for over six months, with tens of thousands of volunteers providing almost four million classifications. After the first third of the images were fully classified, a pre-trained Zoobot was further trained on these and deployed in an active learning cycle with volunteers for the remaining two thirds of the images. A final run of Zoobot was also applied to the entire image data set at the end of the project.
@@
This paper is organised as follows. In Section 2 we discuss the imaging data used for this work, detailing the surveys, subject selection & cutout creation methods, and the metadata associated with each image. In Section 3, further details regarding the project implementation are provided, including the decision tree changes, extra tagging task, the project launch, the use of Zoobot machine learning, and data aggregation at the conclusion of the project. Results are presented in Section 4, including a statistical overview of the various morphological classifications and a comparison between volunteer and Zoobot responses. This is followed by a discussion of the data release catalogues in Section 5 and final conclusions in Section 6.
@@ 2.1 Cosmic Dawn and Hawaii Twenty Square Degree surveys
The Cosmic Dawn survey covers 59 square degrees of the Euclid Deep and Auxiliary Fields from UV/optical to mid-infrared with the aim to understand how galaxies, black holes and dark matter haloes co-evolve from the epoch of reionization to the present – more details of this survey can be found in Euclid Collaboration (2025e). One facet of this is the ongoing H20 survey, which is covering the two primary Euclid deep calibration fields with $u$-band imaging from the CFHT MegaCam instrument and ultra-deep grizgriz imaging from HSC, the latter of which is on board the 8.2-metre Subaru telescope on the summit of Maunakea in Hawaii. H20 aims to push the boundaries of extragalactic astronomy by studying galaxy evolution and co-evolution with dark matter haloes out to a high redshift of $z=7$, through combination with Spitzer Space Telescope infrared imaging and Keck DEIMOS (DEep Imaging Multi-Object Spectrograph) spectroscopy. Targeting the entire twenty square degrees of the EDFN, H20 HSC imaging has intentionally similar depth to Euclid’s own near-infrared imaging in this field, with $5\sigma$ depths of $\sim$27 AB mag (Euclid Collaboration, 2025d) accompanied by a 0.168 arcsec per pixel angular resolution. HSC imaging is seeing-limited, with a median seeing FWHM of 0.67 arcsec over several years of imaging (Miyazaki et al., 2018); variations in the PSF between 0.4 and 1.0 arcsec are observed for the H20 survey (Figure 4 of Euclid Collaboration, 2025d).
The data set used in this iteration of Galaxy Zoo consist of ultra-deep grigri HSC images that cover the first six square degrees of the EDFN observed at full depth by the H20 survey. For reference, the sky coverage of the 9.37 square degrees of the Pre-Launch (PL) EDFN Cosmic Dawn catalogue can be seen in Figure 1 of Euclid Collaboration (2025d), and more details of how the H20 data was obtained and processed can be found in that work.
@@ 2.2 Subject selection
Each subject consists of a deep multi-band cutout image centred on an object identified by the H20 pipeline, limited to those with effective radii $r_{\rm eff}\geq 0.6$ arcsec as the HSC PSF’s 0.67 arcsec FWHM made smaller objects too difficult to classify morphologically. In rare cases (<<1 per cent of all detected sources), the pipeline failed to produce suitable photometric models due to a bright neighbouring object, and so for these sources and point sources the detection-measured major axis was used as a proxy for the effective radius. By taking into account assumptions regarding the average background noise, estimates of the HSC-$i$ band magnitudes for these sources were obtained that were accurate to within $\sim$1 mag, for use in creating image cutouts (see Section 2.3).
@@
Figure 1: Distribution of photometric redshifts for the objects in GZCD as measured by the H20 pipeline. Those for which the pipeline failed to produce suitable photometric redshift estimates are not included. These distributions cover the data set so stars, artifacts and those with ‘bad image zoom’ have not been excluded.
Figure 2: Distribution of HSC-$g$, -$r$, and -$i$ band apparent magnitudes for the objects in GZCD as measured by the H20 pipeline. Those for which the pipeline failed to produce suitable photometric models are not included. These distributions cover the data set so stars, artifacts and those with ‘bad image zoom’ have not been excluded.
The photometric redshift distribution for this subject set can be seen in Figure 1: while 95.4 per cent have $0<z_{\rm phot}<1.0$ (median of $0.42\pm 0.23$), the remaining 4.6 per cent extend up to $z_{\rm phot}<=6.7$. However, it should be noted that for 19.5 per cent of the data set, photometric redshift estimation from the H20 pipeline failed and so these subjects are not included in this range. As discussed in Section 4.1.2, the analysis in later sections is limited to the 45 742 subjects with $z_{\rm phot}<2.5$. Figure 2 shows the distribution of HSC-$g$, -$r$, and -$i$ band apparent magnitudes for the subject set, excluding the small fraction of those for which the modelling failed. Ranges for these magnitudes are $5.64<m_{{\rm HSC}-g}<28.81$, $4.95<m_{{\rm HSC}-r}<29.30$, and $5.22<m_{{\rm HSC}-i}<21.5$, with median averages of 22.03, 21.07, and 20.55, respectively.